jakfishman:

A student is trying to solve the system of two equations given below: Equation P: x + y = 8 Equation Q: 2x + 5y = 24 Which of the following steps can be used to eliminate the x term? (4 points) 2(x + y = 8) −2(x + y = 8) −1(2x + 5y = 24) −2(2x + 5y = 24)

smackz:

ima do the problem partially but then if u can figure out the rest good but if not ill help on it.

smackz:

2x + 2y = 16 (2x + 2y) - (2x + 5y) = 16 - 24 ( modified Equation P minus Equation Q ) 2x - 2x + 2y - 5y = 16 - 24

smackz:

but i beleive first one is your answer.

jhonyy9:

@Laylalyssa any idea here pls ?

jhonyy9:

to eliminate x from the first equation by how many need you multiplie the second equation ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 to eliminate x from the first equation by how many need you multiplie the second equation ? \(\color{#0cbb34}{\text{End of Quote}}\) when add the x terms

jhonyy9:

@XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 to eliminate x from the first equation by how many need you multiplie the second equation ? \(\color{#0cbb34}{\text{End of Quote}}\) when add the x terms \(\color{#0cbb34}{\text{End of Quote}}\) yes and there is no multiplying by -2 but there is multiplying by -4 so -4(a+b=6) is the answer